OpenStudy (unofficialllyy):

Would someone please be willing to check my 9 Geometry questions? I will fan and medal. :)

OpenStudy (unofficialllyy):

1. If a triangle ABC is classified as a right triangle, which statement is true? A. All three angles of triangle ABC are congruent. B. Two slopes of triangle ABC are opposite reciprocals. C. All three sides of triangle ABC are congruent. D. Two slopes of triangle ABC are congruent. I put B.

OpenStudy (unofficialllyy):

@BPDlkeme234 Could you check my answers? :)

OpenStudy (unofficialllyy):

2. Edna has a segment with endpoints F(4, 9) and G(7, 2) that is divided by a point H such that FH and GH form a 1:3 ratio. She knows that the distance between the x-coordinates is 3 units. Which of the following fractions will let her find the x-coordinate for point H? A. 1/4 B. 1/4 C. 3/4 D. 3/1 I put C

OpenStudy (unofficialllyy):

3. Write an equation of a line parallel to line GH below in slope-intercept form that passes through the point (−5, 6). A. y=-x-1 B. y=x-1 C. y=-x+11 D. y=x+11 I put C.